given that A,B,C and D are whole numbers such that A+B=8, B+C=11,B+D=13 and C+D=14, find the values of A,B,C and D.
Found 2 solutions by Cromlix, tommyt3rd:
Answer by Cromlix(4381)

B + C = 11
B + D = 13
Subtract
C - D = -2
C + D = 14
Add
2C = 12
C = 6
B + C = 11
B + 6 = 11
B = 5
A + B = 8
A + 5 = 8
A = 3
B + D = 13
5 + D = 13
D = 8
A = 3, B = 5, C = 6, D = 8
Hope this helps>
:-)

Answer by tommyt3rd(5050)

A+B=8, B+C=11,B+D=13 and C+D=14

(B+C)+(B+D)=11+13
2B+(C+D)=24
2B+14=24
2B=10
B=5

A=3
C=6
D=8

:)